X and Y are continuous R.V. with values X in (0,1), Y in (0,1) and joint PDF f(x,y)=6(x-y)^2, compute the covariance, Cov(X,Y).

Thank you
Answer #1

here E(X)=\int_{0}^{1}\int_{0}^{1} x*f(x,y) dx dy =\int_{0}^{1}\int_{0}^{1} x*(x2+y2-2xy) dx dy =\int_{0}^{1}\int_{0}^{1}(x3+xy2-2x2y) dx dy

=\int_{0}^{1} (x4/4+x2y2/2-2x3y/3) |10 dy =\int_{0}^{1}(1/4+y2/2-2y/3) dy =(y/4+y3/6-y2/3)|10 =(1/4+1/6-1/3)=1/12

similarly E(Y)=\int_{0}^{1}\int_{0}^{1} y*f(x,y) dx dy=1/12

E(XY)=\int_{0}^{1}\int_{0}^{1} xy*f(x,y) dx dy=1/36

hence Cov(X,Y)=E(XY)-E(X)*E(Y)=1/36-(1/12)*(1/12)=3/144=1/48